OpenStudy (faman39):

A truck transports 6 cars to an auto dealership. Each car has a gasoline tank capacity of 64 liters. What is the total gasoline tank capacity of all of the cars in dekaliters? 0.384 dekaliters 3.84 dekaliters 38.4 dekaliters 384 dekaliters

OpenStudy (unklerhaukus):

\[6[\text{cars} ]\times 64[\text L]=6\times 64[\text{cars} ][\text L]\] \[\frac{6\times 64[\text{cars} ][\text L]}{[\text{cars}]}\]\[=6\times 64\frac{[\text{cars} ][\text L]}{[\text{cars}]}\frac{[\text {dL}]}{10[\text L]}\] \[=\frac{6\times 64}{10}\frac{[\text{cars} ][\text L]}{[\text{cars}]}\frac{[\text {dL}]}{[\text L]}\]\[=\]

OpenStudy (faman39):

3 rd one :) thank you so so much @UnkleRhaukus !!!! Your truly really awesome!!! :)

OpenStudy (unklerhaukus):

\[\ddot\smile\]

OpenStudy (faman39):

Thanks alot! :)